# Recursive Training Loops in LLMs: How training data properties modulate distribution shift in generated data?

### v5

@@ C.2.2 Effect on diversity
In section 4.2, we observed that for the Wikipedia dataset, the relationship between diversity loss and synthetic-data ratio was non-linear. Indeed, the greatest drops in diversity are observed for intermediate synthetic-data ratios, rather than for high values as in other datasets and previous works Bertrand et al. (2023); Bohacek and Farid (2023); Kazdan et al. (2024). While at first surprising, we believe this pattern can be explained if we assume that synthetic data aligns more with the models’ priors. This is not a strong assumption given that that data was generated by other fine-tuned versions of the same base models. The intuition, which we experimentally confirm below, is that those datapoints aligned with the model’s priors, have a stronger effect on the training process. And this then leads to the intermediate synthetic data ratios to essentially learn from less data. Let us consider the three different synthetic-data ratios:

### v6

@@ C.2.2 Effect on diversity
In section 4.2, we observed that for the Wikipedia dataset, the relationship between diversity loss and synthetic-data ratio was non-linear. Indeed, the greatest drops in diversity are observed for intermediate synthetic-data ratios, rather than for high values as in other datasets and previous works Bertrand2023; Bohacek2023; Kazdan2024. While at first surprising, we believe this pattern can be explained if we assume that synthetic data aligns more with the models’ priors. This is not a strong assumption given that that data was generated by other fine-tuned versions of the same base models. The intuition, which we experimentally confirm below, is that those datapoints aligned with the model’s priors, have a stronger effect on the training process. And this then leads to the intermediate synthetic data ratios to essentially learn from less data. Let us consider the three different synthetic-data ratios: